You are planning the new layout for the local branch of the Sixth Ninth Bank. You are considering separate cashier windows for the three different classes of service. Each class of service would be separate with its own cashiers and customers. Oddly​ enough, each class of​ service, while​ different, has exactly the same demand and service times. People for one class of service arrive every 3 minutes and arrival times are exponentially distributed​ (the standard deviation is equal to the​ mean). It takes 12 minutes to service each​ customer, and the standard deviation of the service times is 6 minutes. You assign 6 cashiers to each type of service.

a) On​ average, how long will each waiting line be at each of the cashier​ windows?

b) On average how long​ (in minutes) will a customer spend in the​ bank? Assume they​ enter, go directly to one​ line, and leave as soon as service is complete.
Answer #1

The number of cashier. m= 6

Activity time, p = 12 minutes

Arrival time, a = 3 minutes

Utilization , u = p / (a*m) = 12 / (3*6) = 0.6667

CVp = Coefficient of variation of service times = Std deviation of service time / Service time = 6/12 = 0.50

Similarly , CVa = 3/3 =1

Avg length of queue = Lq = ((utiltization^(sqrt(2(m+1)) / (1 - utilization)) * ((CVa^2 + CVp^2)/2) = =(0.6667 ^ SQRT(14)) / (1-0.6667)*(1^2 + 0.5^2)/2 = 0.411

Total in system, Ls = Lq + Number of servers * Utilizaton = 0.411 + 6*0.6667 = 4.411

Average time in queue = Total in system (Ls) / (1/a) = 4.411 / (1/3) =13.233 minutes